LISTSERV mailing list manager LISTSERV 16.0

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


January 2008


View:

Show Author | Hide Author
Show Table of Contents | Hide Table of Contents

 

Table of Contents:

HEA-NOC/20070123-5 [CLOSE] Video conferencing gatekeeper routing rejecting direct IP calls. (1 message)
HEA-NOC/20070813-3 [SUPERCEDE] Repatching Redudant link (1 message)
HEA-NOC/20070918-5 [CLOSE] Repeated scheduled drop from Gi2/0/21 on blanch-colo1 (1 message)
HEA-NOC/20071116-3 [CLOSE] DEG Monitoring Network upgrade (1 message)
HEA-NOC/20071126-7 [CLOSE] MNS services connecting to new network (1 message)
HEA-NOC/20071130-4 [CLOSE] Move INEX Lan #1 link to new router (1 message)
HEA-NOC/20071204-1 [CLOSE] HEAnet Network freeze (18th Dec ~ 7th Jan) (1 message)
HEA-NOC/20071206-6 [SUPERCEDE] LISTSERV upgrade. (1 message)
HEA-NOC/20071206-6 [UPDATE] LISTSERV upgrade. (1 message)
HEA-NOC/20071210-8 [CLOSE] IPv6 tunnel endpoint move (1 message)
HEA-NOC/20071211-6 [CLOSE] Tipp Inst loss of diversity (1 message)
HEA-NOC/20071218-3 [CLOSE] Unscheduled outage on NCRI link (1 message)
HEA-NOC/20071218-8 [CLOSE] Outage in Cork School of Music (1 message)
HEA-NOC/20071219-2 [CLOSE] Kernel updates on multiple machines (1 message)
HEA-NOC/20071219-8 [CLOSE] Electrical works in DKIT (1 message)
HEA-NOC/20071219-8 [UPDATE] Electrical works in DKIT (1 message)
HEA-NOC/20071220-5 [CLOSE] BCFE loss of connectivity/flapping (1 message)
HEA-NOC/20071221-4 [CLOSE] Loss of connectivity to HEA (1 message)
HEA-NOC/20071228-7 [CLOSE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap. (1 message)
HEA-NOC/20071228-7 [UPDATE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap. (1 message)
HEA-NOC/20071231-3 [CLOSE] v6 Routing Issues to a & b.gtld-servers.net (1 message)
HEA-NOC/20080102-1 [CLOSE] Galway PoP Site Failure - Suspected Power Issue (1 message)
HEA-NOC/20080102-1 [OPEN] Galway PoP Site Failure - Suspected Power Issue (1 message)
HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue (4 messages)
HEA-NOC/20080102-6 [CLOSE] IT Carlow UPS Testing (1 message)
HEA-NOC/20080102-6 [OPEN] IT Carlow UPS Testing (1 message)
HEA-NOC/20080102-6 [UPDATE] IT Carlow UPS Testing (1 message)
HEA-NOC/20080103-5 [CLOSE] NUIG Pix firewall outage (1 message)
HEA-NOC/20080103-5 [OPEN] NUIG Pix firewall outage (1 message)
HEA-NOC/20080104-2 [CLOSE] Letterkenny Elastic store overflow at receiver - ADVA (1 message)
HEA-NOC/20080104-2 [OPEN] Letterkenny Elastic store overflow at receiver - ADVA (1 message)
HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA (4 messages)
HEA-NOC/20080104-4 [CLOSE] Citywest services switch reload. (1 message)
HEA-NOC/20080104-4 [OPEN] Citywest services switch reload. (1 message)
HEA-NOC/20080104-6 [CLOSE] GMIT: An Cheim Failover Testing (2 messages)
HEA-NOC/20080104-6 [OPEN] GMIT: An Cheim Failover Testing (1 message)
HEA-NOC/20080104-6 [REOPEN] GMIT: An Cheim Failover Testing (1 message)
HEA-NOC/20080107-1 [CLOSE] Bluenet Interfaces to be changed from Routing to Switching mode. (1 message)
HEA-NOC/20080107-1 [OPEN] Bluenet Interfaces to be changed from Routing to Switching mode. (1 message)
HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode. (8 messages)
HEA-NOC/20080107-4 [CLOSE] HEAnet Network Freeze (17th ~ 21st Jan 2008) (1 message)
HEA-NOC/20080107-4 [OPEN] HEAnet Network Freeze (17th ~ 21st Jan 2008) (1 message)
HEA-NOC/20080107-6 [OPEN] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008) (1 message)
HEA-NOC/20080107-6 [UPDATE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008) (1 message)
HEA-NOC/20080107-8 [CLOSE] IADT Dun Laoghaire - HEAnet Migration (1 message)
HEA-NOC/20080107-8 [OPEN] IADT Dun Laoghaire - HEAnet Migration (1 message)
HEA-NOC/20080107-8 [UPDATE] IADT Dun Laoghaire - HEAnet Migration (1 message)
HEA-NOC/20080109-1 [CLOSE] CIT An Chem failover testing (1 message)
HEA-NOC/20080109-1 [OPEN] CIT An Chem failover testing (1 message)
HEA-NOC/20080109-1 [UPDATE] CIT An Chem failover testing (1 message)
HEA-NOC/20080109-10 [CLOSE] Dundalk power outage and loss of resilience to Galway sites (1 message)
HEA-NOC/20080109-10 [OPEN] Dundalk power outage and loss of resilience to Galway sites (1 message)
HEA-NOC/20080109-10 [UPDATE] Dundalk power outage and loss of resilience to Galway sites (2 messages)
HEA-NOC/20080109-7 [CLOSE] An Cheim interruption (1 message)
HEA-NOC/20080109-7 [OPEN] An Cheim failover (1 message)
HEA-NOC/20080109-7 [UPDATE] An Cheim failover (2 messages)
HEA-NOC/20080109-7 [UPDATE] An Cheim interruption (1 message)
HEA-NOC/20080110-1 [CLOSE] Dun Laoghaire IADT: An Cheim Failover Testing (1 message)
HEA-NOC/20080110-1 [OPEN] Dun Laoghaire IADT: An Cheim Failover Testing (1 message)
HEA-NOC/20080110-1 [UPDATE] Dun Laoghaire IADT: An Cheim Failover Testing (3 messages)
HEA-NOC/20080111-11 [OPEN/CLOSE] Power Outage in Cork (1 message)
HEA-NOC/20080111-13 [CLOSE] Loss of resilience on Dublin core ring (1 message)
HEA-NOC/20080111-13 [OPEN] Loss of resilience on Dublin core ring (1 message)
HEA-NOC/20080111-13 [UPDATE] Loss of resilience on Dublin core ring (2 messages)
HEA-NOC/20080111-4 [CLOSE] Drupal upgrade required (1 message)
HEA-NOC/20080111-4 [OPEN] Drupal upgrade required (1 message)
HEA-NOC/20080111-6 [CLOSE] HEAnet Co-Lo KVM Service Inoperable (1 message)
HEA-NOC/20080111-6 [OPEN] HEAnet Co-Lo KVM Service Inoperable (1 message)
HEA-NOC/20080111-9 [CLOSE] Flaps to VECs Crumlin, Killester and Liberties (1 message)
HEA-NOC/20080111-9 [OPEN] Flaps to VECs Crumlin, Killester and Liberties (1 message)
HEA-NOC/20080111-9 [UPDATE] Flaps to VECs Crumlin, Killester and Liberties (2 messages)
HEA-NOC/20080112-1 [OPEN] Multiple flaps on connecting circuit to BFCE (1 message)
HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE (8 messages)
HEA-NOC/20080114-4 [OPEN/CLOSE] Power Outage in Cork PoP (1 message)
HEA-NOC/20080114-6 [CLOSE] Wordpress upgrade required (1 message)
HEA-NOC/20080114-6 [OPEN] Wordpress upgrade required (1 message)
HEA-NOC/20080116-13 [OPEN] Establish peerings between NCI & the HEAnet new core routers (1 message)
HEA-NOC/20080116-13 [UPDATE] Establish peerings between NCI & the HEAnet new core routers (2 messages)
HEA-NOC/20080116-2 [CLOSE] Link to GEANT down. (1 message)
HEA-NOC/20080116-2 [OPEN] Link to GEANT down. (1 message)
HEA-NOC/20080116-2 [UPDATE] Link to GEANT down. (3 messages)
HEA-NOC/20080116-7 [CLOSE] BT fibre maintanence on Cork-Limerick GigE link (1 message)
HEA-NOC/20080116-7 [OPEN] BT fibre maintanence on Cork-Limerick GigE link (1 message)
HEA-NOC/20080118-3 [OPEN] Both circuits to the IEDR are reporting down (1 message)
HEA-NOC/20080118-3 [UPDATE] Both circuits to the IEDR are reporting down (1 message)
HEA-NOC/20080118-4 [CLOSE] UL website not responding (1 message)
HEA-NOC/20080118-4 [OPEN] UL website not responding (1 message)
HEA-NOC/20080118-4 [UPDATE] UL website not responding (1 message)
HEA-NOC/20080119-2 [CLOSE] GMIT inaccessible (1 message)
HEA-NOC/20080119-2 [OPEN] GMIT inaccessible (1 message)
HEA-NOC/20080119-2 [UPDATE] GMIT inaccessible (1 message)
HEA-NOC/20080121-12 [OPEN] ESBT Planned Outage Notice (1 message)
HEA-NOC/20080121-2 [CLOSE] ESB Outage in Dundalk (1 message)
HEA-NOC/20080121-2 [OPEN] ESB Outage in Dundalk (1 message)
HEA-NOC/20080121-2 [UPDATE] ESB Outage in Dundalk (1 message)
HEA-NOC/20080121-9 [OPEN] IOS Upgrades (1 message)
HEA-NOC/20080122-5 [CLOSE] Maintenance Work CDVEC (1 message)
HEA-NOC/20080122-5 [OPEN] Maintenance Work CDVEC (1 message)
HEA-NOC/20080122-6 [CLOSE] DIT migration to new IP equipment (1 message)
HEA-NOC/20080122-6 [OPEN] DIT migration to new IP equipment (1 message)
HEA-NOC/20080122-9 [CLOSE] EPA Firewall Testing (1 message)
HEA-NOC/20080122-9 [OPEN] EPA Firewall TestingThe (1 message)
HEA-NOC/20080122-9 [UPDATE] EPA Firewall Testing (2 messages)
HEA-NOC/20080122-9 [UPDATE] EPA Firewall TestingThe (1 message)
HEA-NOC/20080123-10 [CLOSE] mirror services degraded. (1 message)
HEA-NOC/20080123-10 [OPEN] mirror services degraded. (1 message)
HEA-NOC/20080123-10 [UPDATE] mirror services degraded. (4 messages)
HEA-NOC/20080123-11 [CLOSE] Migration of LISTSERV to new server. (1 message)
HEA-NOC/20080123-11 [OPEN] Migration of LISTSERV to new server. (1 message)
HEA-NOC/20080123-11 [UPDATE] Migration of LISTSERV to new server. (3 messages)
HEA-NOC/20080123-12 [CLOSE] Change of Bluenet PE node interfaces from routing to switching (1 message)
HEA-NOC/20080123-12 [OPEN] Change of Bluenet PE node interfaces from routing to switching (1 message)
HEA-NOC/20080123-12 [UPDATE] Change of Bluenet PE node interfaces from routing to switching (3 messages)
HEA-NOC/20080123-8 [CLOSE] Loss of connectivity (1 message)
HEA-NOC/20080123-8 [OPEN] Loss of connectivity (1 message)
HEA-NOC/20080123-8 [UPDATE] Loss of connectivity (2 messages)
HEA-NOC/20080124-1 [OPEN] DIAS IOS upgrade. (1 message)
HEA-NOC/20080124-1 [UPDATE] DIAS IOS upgrade. (1 message)
HEA-NOC/20080124-4 [OPEN] Janet connection down (1 message)
HEA-NOC/20080124-4 [UPDATE] Janet connection down (1 message)
HEA-NOC/20080124-5 [CLOSE] Loss of connectivity Outreach Centre UL (1 message)
HEA-NOC/20080124-5 [OPEN] Loss of connectivity Outreach Centre UL (1 message)
HEA-NOC/20080125-1 [CLOSE] Loss of connectivity (1 message)
HEA-NOC/20080125-1 [OPEN] Loss of connectivity (1 message)
HEA-NOC/20080125-1 [UPDATE] Loss of connectivity (4 messages)
HEA-NOC/20080125-3 [CLOSE] Upgrade coupland to FreeBSD 6.3 (1 message)
HEA-NOC/20080125-3 [OPEN] Upgrade coupland to FreeBSD 6.3 (1 message)
HEA-NOC/20080126-1 [CLOSE] Hermes intermittent outages (1 message)
HEA-NOC/20080126-1 [OPEN] Hermes intermittent outages (1 message)
HEA-NOC/20080126-1 [UPDATE] Hermes intermittent outages (2 messages)
HEA-NOC/20080126-5 [OPEN] Bluenet Dublin ring brief loss of resilience (1 message)
HEA-NOC/20080127-1 [CLOSE] GEANT primary circuit outage (1 message)
HEA-NOC/20080127-1 [OPEN] GEANT primary circuit outage (1 message)
HEA-NOC/20080127-1 [UPDATE] GEANT primary circuit outage (1 message)
HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap (1 message)
HEA-NOC/20080127-5 [OPEN] St. Angela's brief flap (1 message)
HEA-NOC/20080127-5 [REOPEN] St. Angela's brief flap (1 message)
HEA-NOC/20080127-5 [UPDATE] St. Angela's brief flap (3 messages)
HEA-NOC/20080128-2 [CLOSE] IT Tallaght migration to new IP equipment (1 message)
HEA-NOC/20080128-2 [UPDATE] IT Tallaght migration to new IP equipment (1 message)
HEA-NOC/20080129-10 [OPEN] Upgrade to SNMP module on centralised DC Rectifier (1 message)
HEA-NOC/20080129-2 [CLOSE] Outage on primary GEANT link (1 message)
HEA-NOC/20080129-2 [OPEN] Outage on primary GEANT link (1 message)
HEA-NOC/20080129-9 [OPEN] WIT migration to new IP equipment (1 message)
HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment (1 message)
HEA-NOC/20080130-2 [CLOSE] Loss of connectivity to parts of Services network in Blanch (1 message)
HEA-NOC/20080130-2 [OPEN] Loss of connectivity to parts of Services network in Blanch (1 message)
HEA-NOC/20080130-2 [UPDATE] Loss of connectivity to parts of Services network in Blanch (1 message)

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

Subject Sorted by Subject, Chronologically

From

Date

Size

New Thread

HEA-NOC/20070123-5 [CLOSE] Video conferencing gatekeeper routing rejecting direct IP calls.

HEA-NOC/20070123-5 [CLOSE] Video conferencing gatekeeper routing rejecting direct IP calls.

Justin Hourigan <[log in to unmask]>

Tue, 8 Jan 2008 11:43:17 +0000

55 lines

New Thread

HEA-NOC/20070813-3 [SUPERCEDE] Repatching Redudant link

HEA-NOC/20070813-3 [SUPERCEDE] Repatching Redudant link

Gareth Eason <[log in to unmask]>

Fri, 4 Jan 2008 17:08:32 +0000

40 lines

New Thread

HEA-NOC/20070918-5 [CLOSE] Repeated scheduled drop from Gi2/0/21 on blanch-colo1

HEA-NOC/20070918-5 [CLOSE] Repeated scheduled drop from Gi2/0/21 on blanch-colo1

Gareth Eason <[log in to unmask]>

Sun, 13 Jan 2008 10:47:30 +0000

59 lines

New Thread

HEA-NOC/20071116-3 [CLOSE] DEG Monitoring Network upgrade

HEA-NOC/20071116-3 [CLOSE] DEG Monitoring Network upgrade

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 12:55:56 +0000

30 lines

New Thread

HEA-NOC/20071126-7 [CLOSE] MNS services connecting to new network

HEA-NOC/20071126-7 [CLOSE] MNS services connecting to new network

Dave Wilson <[log in to unmask]>

Mon, 7 Jan 2008 13:54:47 +0000

68 lines

New Thread

HEA-NOC/20071130-4 [CLOSE] Move INEX Lan #1 link to new router

HEA-NOC/20071130-4 [CLOSE] Move INEX Lan #1 link to new router

Dave Wilson <[log in to unmask]>

Mon, 7 Jan 2008 13:55:33 +0000

52 lines

New Thread

HEA-NOC/20071204-1 [CLOSE] HEAnet Network freeze (18th Dec ~ 7th Jan)

HEA-NOC/20071204-1 [CLOSE] HEAnet Network freeze (18th Dec ~ 7th Jan)

Gareth Eason <[log in to unmask]>

Tue, 8 Jan 2008 17:59:10 +0000

70 lines

New Thread

HEA-NOC/20071206-6 [SUPERCEDE] LISTSERV upgrade.

HEA-NOC/20071206-6 [SUPERCEDE] LISTSERV upgrade.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 16:28:31 +0000

40 lines

New Thread

HEA-NOC/20071206-6 [UPDATE] LISTSERV upgrade.

HEA-NOC/20071206-6 [UPDATE] LISTSERV upgrade.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 15:41:16 +0000

37 lines

New Thread

HEA-NOC/20071210-8 [CLOSE] IPv6 tunnel endpoint move

HEA-NOC/20071210-8 [CLOSE] IPv6 tunnel endpoint move

Dave Wilson <[log in to unmask]>

Mon, 7 Jan 2008 13:59:10 +0000

53 lines

New Thread

HEA-NOC/20071211-6 [CLOSE] Tipp Inst loss of diversity

HEA-NOC/20071211-6 [CLOSE] Tipp Inst loss of diversity

Dave Wilson <[log in to unmask]>

Mon, 7 Jan 2008 13:58:21 +0000

37 lines

New Thread

HEA-NOC/20071218-3 [CLOSE] Unscheduled outage on NCRI link

HEA-NOC/20071218-3 [CLOSE] Unscheduled outage on NCRI link

Rob Gallagher <[log in to unmask]>

Tue, 8 Jan 2008 14:31:08 +0000

37 lines

New Thread

HEA-NOC/20071218-8 [CLOSE] Outage in Cork School of Music

HEA-NOC/20071218-8 [CLOSE] Outage in Cork School of Music

Rob Gallagher <[log in to unmask]>

Tue, 8 Jan 2008 14:32:12 +0000

36 lines

New Thread

HEA-NOC/20071219-2 [CLOSE] Kernel updates on multiple machines

HEA-NOC/20071219-2 [CLOSE] Kernel updates on multiple machines

Gareth Eason <[log in to unmask]>

Wed, 9 Jan 2008 17:30:06 +0000

66 lines

New Thread

HEA-NOC/20071219-8 [CLOSE] Electrical works in DKIT

HEA-NOC/20071219-8 [CLOSE] Electrical works in DKIT

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 16:09:26 +0000

44 lines

New Thread

HEA-NOC/20071219-8 [UPDATE] Electrical works in DKIT

HEA-NOC/20071219-8 [UPDATE] Electrical works in DKIT

Garwin Liu <[log in to unmask]>

Fri, 4 Jan 2008 11:58:40 +0000

41 lines

New Thread

HEA-NOC/20071220-5 [CLOSE] BCFE loss of connectivity/flapping

HEA-NOC/20071220-5 [CLOSE] BCFE loss of connectivity/flapping

Rob Gallagher <[log in to unmask]>

Wed, 2 Jan 2008 12:14:49 +0000

35 lines

New Thread

HEA-NOC/20071221-4 [CLOSE] Loss of connectivity to HEA

HEA-NOC/20071221-4 [CLOSE] Loss of connectivity to HEA

Rob Gallagher <[log in to unmask]>

Wed, 2 Jan 2008 12:13:30 +0000

31 lines

New Thread

HEA-NOC/20071228-7 [CLOSE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap.

HEA-NOC/20071228-7 [CLOSE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap.

Brian Scanlan <[log in to unmask]>

Mon, 7 Jan 2008 11:33:15 +0000

44 lines

New Thread

HEA-NOC/20071228-7 [UPDATE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap.

HEA-NOC/20071228-7 [UPDATE] HEAnet - Geant Secondary to Geneva (Bluenet ID: 274) flap.

Brian Scanlan <[log in to unmask]>

Wed, 2 Jan 2008 14:12:29 +0000

42 lines

New Thread

HEA-NOC/20071231-3 [CLOSE] v6 Routing Issues to a & b.gtld-servers.net

HEA-NOC/20071231-3 [CLOSE] v6 Routing Issues to a & b.gtld-servers.net

Brian Nisbet <[log in to unmask]>

Wed, 2 Jan 2008 09:34:52 +0000

44 lines

New Thread

HEA-NOC/20080102-1 [CLOSE] Galway PoP Site Failure - Suspected Power Issue

HEA-NOC/20080102-1 [CLOSE] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Fri, 4 Jan 2008 10:07:36 +0000

52 lines

New Thread

HEA-NOC/20080102-1 [OPEN] Galway PoP Site Failure - Suspected Power Issue

HEA-NOC/20080102-1 [OPEN] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 09:03:17 +0000

32 lines

New Thread

HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue

HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 11:59:42 +0000

48 lines

HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 11:28:49 +0000

43 lines

HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 10:24:03 +0000

41 lines

HEA-NOC/20080102-1 [UPDATE] Galway PoP Site Failure - Suspected Power Issue

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 09:28:35 +0000

38 lines

New Thread

HEA-NOC/20080102-6 [CLOSE] IT Carlow UPS Testing

HEA-NOC/20080102-6 [CLOSE] IT Carlow UPS Testing

Garwin Liu <[log in to unmask]>

Fri, 4 Jan 2008 10:05:13 +0000

37 lines

New Thread

HEA-NOC/20080102-6 [OPEN] IT Carlow UPS Testing

HEA-NOC/20080102-6 [OPEN] IT Carlow UPS Testing

Garwin Liu <[log in to unmask]>

Wed, 2 Jan 2008 15:54:01 +0000

29 lines

New Thread

HEA-NOC/20080102-6 [UPDATE] IT Carlow UPS Testing

HEA-NOC/20080102-6 [UPDATE] IT Carlow UPS Testing

Garwin Liu <[log in to unmask]>

Thu, 3 Jan 2008 10:16:00 +0000

32 lines

New Thread

HEA-NOC/20080103-5 [CLOSE] NUIG Pix firewall outage

HEA-NOC/20080103-5 [CLOSE] NUIG Pix firewall outage

Garreth Malone <[log in to unmask]>

Sat, 5 Jan 2008 15:43:23 +0000

40 lines

New Thread

HEA-NOC/20080103-5 [OPEN] NUIG Pix firewall outage

HEA-NOC/20080103-5 [OPEN] NUIG Pix firewall outage

Garwin Liu <[log in to unmask]>

Thu, 3 Jan 2008 13:30:43 +0000

36 lines

New Thread

HEA-NOC/20080104-2 [CLOSE] Letterkenny Elastic store overflow at receiver - ADVA

HEA-NOC/20080104-2 [CLOSE] Letterkenny Elastic store overflow at receiver - ADVA

Garreth Malone <[log in to unmask]>

Mon, 7 Jan 2008 11:54:34 +0000

58 lines

New Thread

HEA-NOC/20080104-2 [OPEN] Letterkenny Elastic store overflow at receiver - ADVA

HEA-NOC/20080104-2 [OPEN] Letterkenny Elastic store overflow at receiver - ADVA

Garwin Liu <[log in to unmask]>

Fri, 4 Jan 2008 09:31:05 +0000

27 lines

New Thread

HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA

HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA

Garreth Malone <[log in to unmask]>

Sat, 5 Jan 2008 12:54:52 +0000

51 lines

HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA

Garwin Liu <[log in to unmask]>

Fri, 4 Jan 2008 16:15:28 +0000

48 lines

HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA

Gareth Eason <[log in to unmask]>

Fri, 4 Jan 2008 12:12:07 +0000

39 lines

HEA-NOC/20080104-2 [UPDATE] Letterkenny Elastic store overflow at receiver - ADVA

Gareth Eason <[log in to unmask]>

Fri, 4 Jan 2008 10:19:57 +0000

35 lines

New Thread

HEA-NOC/20080104-4 [CLOSE] Citywest services switch reload.

HEA-NOC/20080104-4 [CLOSE] Citywest services switch reload.

Brian Scanlan <[log in to unmask]>

Wed, 16 Jan 2008 09:28:54 +0000

34 lines

New Thread

HEA-NOC/20080104-4 [OPEN] Citywest services switch reload.

HEA-NOC/20080104-4 [OPEN] Citywest services switch reload.

Brian Scanlan <[log in to unmask]>

Fri, 4 Jan 2008 11:54:34 +0000

28 lines

New Thread

HEA-NOC/20080104-6 [CLOSE] GMIT: An Cheim Failover Testing

HEA-NOC/20080104-6 [CLOSE] GMIT: An Cheim Failover Testing

Owen Byrne <[log in to unmask]>

Tue, 8 Jan 2008 14:13:12 +0000

45 lines

HEA-NOC/20080104-6 [CLOSE] GMIT: An Cheim Failover Testing

Owen Byrne <[log in to unmask]>

Tue, 8 Jan 2008 09:47:34 +0000

40 lines

New Thread

HEA-NOC/20080104-6 [OPEN] GMIT: An Cheim Failover Testing

HEA-NOC/20080104-6 [OPEN] GMIT: An Cheim Failover Testing

Owen Byrne <[log in to unmask]>

Fri, 4 Jan 2008 13:03:01 +0000

34 lines

New Thread

HEA-NOC/20080104-6 [REOPEN] GMIT: An Cheim Failover Testing

HEA-NOC/20080104-6 [REOPEN] GMIT: An Cheim Failover Testing

Owen Byrne <[log in to unmask]>

Tue, 8 Jan 2008 14:12:20 +0000

40 lines

New Thread

HEA-NOC/20080107-1 [CLOSE] Bluenet Interfaces to be changed from Routing to Switching mode.

HEA-NOC/20080107-1 [CLOSE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Mon, 21 Jan 2008 17:38:59 +0000

100 lines

New Thread

HEA-NOC/20080107-1 [OPEN] Bluenet Interfaces to be changed from Routing to Switching mode.

HEA-NOC/20080107-1 [OPEN] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Mon, 7 Jan 2008 10:24:14 +0000

49 lines

New Thread

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 16 Jan 2008 09:08:23 +0000

91 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 16 Jan 2008 08:17:16 +0000

85 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 16 Jan 2008 07:11:55 +0000

79 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Tue, 15 Jan 2008 13:07:37 +0000

75 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 9 Jan 2008 20:56:57 +0000

68 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 9 Jan 2008 09:22:12 +0000

61 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 9 Jan 2008 09:05:16 +0000

58 lines

HEA-NOC/20080107-1 [UPDATE] Bluenet Interfaces to be changed from Routing to Switching mode.

Gareth Eason <[log in to unmask]>

Wed, 9 Jan 2008 07:20:30 +0000

54 lines

New Thread

HEA-NOC/20080107-4 [CLOSE] HEAnet Network Freeze (17th ~ 21st Jan 2008)

HEA-NOC/20080107-4 [CLOSE] HEAnet Network Freeze (17th ~ 21st Jan 2008)

Gareth Eason <[log in to unmask]>

Mon, 21 Jan 2008 17:22:15 +0000

50 lines

New Thread

HEA-NOC/20080107-4 [OPEN] HEAnet Network Freeze (17th ~ 21st Jan 2008)

HEA-NOC/20080107-4 [OPEN] HEAnet Network Freeze (17th ~ 21st Jan 2008)

Gareth Eason <[log in to unmask]>

Mon, 7 Jan 2008 11:15:08 +0000

47 lines

New Thread

HEA-NOC/20080107-6 [OPEN] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

HEA-NOC/20080107-6 [OPEN] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

Gareth Eason <[log in to unmask]>

Mon, 7 Jan 2008 11:18:34 +0000

49 lines

New Thread

HEA-NOC/20080107-6 [UPDATE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

HEA-NOC/20080107-6 [UPDATE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

Gareth Eason <[log in to unmask]>

Thu, 31 Jan 2008 10:05:23 +0000

53 lines

New Thread

HEA-NOC/20080107-8 [CLOSE] IADT Dun Laoghaire - HEAnet Migration

HEA-NOC/20080107-8 [CLOSE] IADT Dun Laoghaire - HEAnet Migration

Owen Byrne <[log in to unmask]>

Wed, 9 Jan 2008 11:10:15 +0000

44 lines

New Thread

HEA-NOC/20080107-8 [OPEN] IADT Dun Laoghaire - HEAnet Migration

HEA-NOC/20080107-8 [OPEN] IADT Dun Laoghaire - HEAnet Migration

Owen Byrne <[log in to unmask]>

Mon, 7 Jan 2008 14:31:48 +0000

31 lines

New Thread

HEA-NOC/20080107-8 [UPDATE] IADT Dun Laoghaire - HEAnet Migration

HEA-NOC/20080107-8 [UPDATE] IADT Dun Laoghaire - HEAnet Migration

Garwin Liu <[log in to unmask]>

Tue, 8 Jan 2008 16:13:38 +0000

36 lines

New Thread

HEA-NOC/20080109-1 [CLOSE] CIT An Chem failover testing

HEA-NOC/20080109-1 [CLOSE] CIT An Chem failover testing

Garwin Liu <[log in to unmask]>

Thu, 17 Jan 2008 15:52:26 +0000

43 lines

New Thread

HEA-NOC/20080109-1 [OPEN] CIT An Chem failover testing

HEA-NOC/20080109-1 [OPEN] CIT An Chem failover testing

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 10:26:09 +0000

35 lines

New Thread

HEA-NOC/20080109-1 [UPDATE] CIT An Chem failover testing

HEA-NOC/20080109-1 [UPDATE] CIT An Chem failover testing

Garwin Liu <[log in to unmask]>

Tue, 15 Jan 2008 12:38:49 +0000

39 lines

New Thread

HEA-NOC/20080109-10 [CLOSE] Dundalk power outage and loss of resilience to Galway sites

HEA-NOC/20080109-10 [CLOSE] Dundalk power outage and loss of resilience to Galway sites

Dave Wilson <[log in to unmask]>

Mon, 14 Jan 2008 11:39:30 +0000

46 lines

New Thread

HEA-NOC/20080109-10 [OPEN] Dundalk power outage and loss of resilience to Galway sites

HEA-NOC/20080109-10 [OPEN] Dundalk power outage and loss of resilience to Galway sites

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 16:35:30 +0000

35 lines

New Thread

HEA-NOC/20080109-10 [UPDATE] Dundalk power outage and loss of resilience to Galway sites

HEA-NOC/20080109-10 [UPDATE] Dundalk power outage and loss of resilience to Galway sites

Dave Wilson <[log in to unmask]>

Fri, 11 Jan 2008 17:23:23 +0000

43 lines

HEA-NOC/20080109-10 [UPDATE] Dundalk power outage and loss of resilience to Galway sites

Dave Wilson <[log in to unmask]>

Fri, 11 Jan 2008 16:49:27 +0000

39 lines

New Thread

HEA-NOC/20080109-7 [CLOSE] An Cheim interruption

HEA-NOC/20080109-7 [CLOSE] An Cheim interruption

Dave Wilson <[log in to unmask]>

Mon, 21 Jan 2008 11:17:53 +0000

55 lines

New Thread

HEA-NOC/20080109-7 [OPEN] An Cheim failover

HEA-NOC/20080109-7 [OPEN] An Cheim failover

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 14:37:40 +0000

33 lines

New Thread

HEA-NOC/20080109-7 [UPDATE] An Cheim failover

HEA-NOC/20080109-7 [UPDATE] An Cheim failover

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 15:01:53 +0000

49 lines

HEA-NOC/20080109-7 [UPDATE] An Cheim failover

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 14:54:52 +0000

46 lines

New Thread

HEA-NOC/20080109-7 [UPDATE] An Cheim interruption

HEA-NOC/20080109-7 [UPDATE] An Cheim interruption

Dave Wilson <[log in to unmask]>

Wed, 9 Jan 2008 15:10:07 +0000

53 lines

New Thread

HEA-NOC/20080110-1 [CLOSE] Dun Laoghaire IADT: An Cheim Failover Testing

HEA-NOC/20080110-1 [CLOSE] Dun Laoghaire IADT: An Cheim Failover Testing

Garwin Liu <[log in to unmask]>

Mon, 14 Jan 2008 13:17:50 +0000

48 lines

New Thread

HEA-NOC/20080110-1 [OPEN] Dun Laoghaire IADT: An Cheim Failover Testing

HEA-NOC/20080110-1 [OPEN] Dun Laoghaire IADT: An Cheim Failover Testing

Garwin Liu <[log in to unmask]>

Thu, 10 Jan 2008 09:20:53 +0000

31 lines

New Thread

HEA-NOC/20080110-1 [UPDATE] Dun Laoghaire IADT: An Cheim Failover Testing

HEA-NOC/20080110-1 [UPDATE] Dun Laoghaire IADT: An Cheim Failover Testing

Garwin Liu <[log in to unmask]>

Mon, 14 Jan 2008 13:17:31 +0000

46 lines

HEA-NOC/20080110-1 [UPDATE] Dun Laoghaire IADT: An Cheim Failover Testing

Garwin Liu <[log in to unmask]>

Fri, 11 Jan 2008 15:05:31 +0000

39 lines

HEA-NOC/20080110-1 [UPDATE] Dun Laoghaire IADT: An Cheim Failover Testing

Garwin Liu <[log in to unmask]>

Thu, 10 Jan 2008 13:39:55 +0000

37 lines

New Thread

HEA-NOC/20080111-11 [OPEN/CLOSE] Power Outage in Cork

HEA-NOC/20080111-11 [OPEN/CLOSE] Power Outage in Cork

Brian Nisbet <[log in to unmask]>

Fri, 11 Jan 2008 12:54:26 +0000

33 lines

New Thread

HEA-NOC/20080111-13 [CLOSE] Loss of resilience on Dublin core ring

HEA-NOC/20080111-13 [CLOSE] Loss of resilience on Dublin core ring

Dave Wilson <[log in to unmask]>

Mon, 21 Jan 2008 11:17:21 +0000

45 lines

New Thread

HEA-NOC/20080111-13 [OPEN] Loss of resilience on Dublin core ring

HEA-NOC/20080111-13 [OPEN] Loss of resilience on Dublin core ring

Dave Wilson <[log in to unmask]>

Fri, 11 Jan 2008 15:28:55 +0000

36 lines

New Thread

HEA-NOC/20080111-13 [UPDATE] Loss of resilience on Dublin core ring

HEA-NOC/20080111-13 [UPDATE] Loss of resilience on Dublin core ring

Dave Wilson <[log in to unmask]>

Mon, 14 Jan 2008 10:24:51 +0000

44 lines

HEA-NOC/20080111-13 [UPDATE] Loss of resilience on Dublin core ring

Justin Hourigan <[log in to unmask]>

Mon, 14 Jan 2008 09:21:36 +0000

41 lines

New Thread

HEA-NOC/20080111-4 [CLOSE] Drupal upgrade required

HEA-NOC/20080111-4 [CLOSE] Drupal upgrade required

Rob Gallagher <[log in to unmask]>

Fri, 11 Jan 2008 15:59:33 +0000

49 lines

New Thread

HEA-NOC/20080111-4 [OPEN] Drupal upgrade required

HEA-NOC/20080111-4 [OPEN] Drupal upgrade required

Rob Gallagher <[log in to unmask]>

Fri, 11 Jan 2008 10:14:46 +0000

43 lines

New Thread

HEA-NOC/20080111-6 [CLOSE] HEAnet Co-Lo KVM Service Inoperable

HEA-NOC/20080111-6 [CLOSE] HEAnet Co-Lo KVM Service Inoperable

Brian Nisbet <[log in to unmask]>

Fri, 11 Jan 2008 10:52:37 +0000

36 lines

New Thread

HEA-NOC/20080111-6 [OPEN] HEAnet Co-Lo KVM Service Inoperable

HEA-NOC/20080111-6 [OPEN] HEAnet Co-Lo KVM Service Inoperable

Brian Nisbet <[log in to unmask]>

Fri, 11 Jan 2008 10:48:31 +0000

32 lines

New Thread

HEA-NOC/20080111-9 [CLOSE] Flaps to VECs Crumlin, Killester and Liberties

HEA-NOC/20080111-9 [CLOSE] Flaps to VECs Crumlin, Killester and Liberties

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 16:04:54 +0000

48 lines

New Thread

HEA-NOC/20080111-9 [OPEN] Flaps to VECs Crumlin, Killester and Liberties

HEA-NOC/20080111-9 [OPEN] Flaps to VECs Crumlin, Killester and Liberties

Dave Wilson <[log in to unmask]>

Fri, 11 Jan 2008 12:34:33 +0000

33 lines

New Thread

HEA-NOC/20080111-9 [UPDATE] Flaps to VECs Crumlin, Killester and Liberties

HEA-NOC/20080111-9 [UPDATE] Flaps to VECs Crumlin, Killester and Liberties

Gareth Eason <[log in to unmask]>

Sat, 12 Jan 2008 15:45:50 +0000

46 lines

HEA-NOC/20080111-9 [UPDATE] Flaps to VECs Crumlin, Killester and Liberties

Dave Wilson <[log in to unmask]>

Fri, 11 Jan 2008 17:07:49 +0000

38 lines

New Thread

HEA-NOC/20080112-1 [OPEN] Multiple flaps on connecting circuit to BFCE

HEA-NOC/20080112-1 [OPEN] Multiple flaps on connecting circuit to BFCE

Gareth Eason <[log in to unmask]>

Sat, 12 Jan 2008 10:48:30 +0000

30 lines

New Thread

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Gareth Eason <[log in to unmask]>

Thu, 24 Jan 2008 11:55:26 +0000

58 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Gareth Eason <[log in to unmask]>

Wed, 23 Jan 2008 14:58:04 +0000

54 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Cormac O'Cianain <[log in to unmask]>

Wed, 23 Jan 2008 13:03:31 +0000

51 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Justin Hourigan <[log in to unmask]>

Thu, 17 Jan 2008 15:15:54 +0000

49 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Justin Hourigan <[log in to unmask]>

Thu, 17 Jan 2008 12:32:32 +0000

46 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Justin Hourigan <[log in to unmask]>

Thu, 17 Jan 2008 12:30:23 +0000

43 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Justin Hourigan <[log in to unmask]>

Mon, 14 Jan 2008 11:18:56 +0000

39 lines

HEA-NOC/20080112-1 [UPDATE] Multiple flaps on connecting circuit to BFCE

Justin Hourigan <[log in to unmask]>

Mon, 14 Jan 2008 10:34:07 +0000

35 lines

New Thread

HEA-NOC/20080114-4 [OPEN/CLOSE] Power Outage in Cork PoP

HEA-NOC/20080114-4 [OPEN/CLOSE] Power Outage in Cork PoP

Brian Nisbet <[log in to unmask]>

Mon, 14 Jan 2008 10:07:46 +0000

35 lines

New Thread

HEA-NOC/20080114-6 [CLOSE] Wordpress upgrade required

HEA-NOC/20080114-6 [CLOSE] Wordpress upgrade required

Rob Gallagher <[log in to unmask]>

Mon, 14 Jan 2008 12:23:54 +0000

47 lines

New Thread

HEA-NOC/20080114-6 [OPEN] Wordpress upgrade required

HEA-NOC/20080114-6 [OPEN] Wordpress upgrade required

Rob Gallagher <[log in to unmask]>

Mon, 14 Jan 2008 12:03:22 +0000

44 lines

New Thread

HEA-NOC/20080116-13 [OPEN] Establish peerings between NCI & the HEAnet new core routers

HEA-NOC/20080116-13 [OPEN] Establish peerings between NCI & the HEAnet new core routers

Paul Cosgrove <[log in to unmask]>

Wed, 16 Jan 2008 17:53:57 +0000

45 lines

New Thread

HEA-NOC/20080116-13 [UPDATE] Establish peerings between NCI & the HEAnet new core routers

HEA-NOC/20080116-13 [UPDATE] Establish peerings between NCI & the HEAnet new core routers

Paul Cosgrove <[log in to unmask]>

Thu, 17 Jan 2008 11:52:07 +0000

58 lines

HEA-NOC/20080116-13 [UPDATE] Establish peerings between NCI & the HEAnet new core routers

Paul Cosgrove <[log in to unmask]>

Thu, 17 Jan 2008 10:51:23 +0000

52 lines

New Thread

HEA-NOC/20080116-2 [CLOSE] Link to GEANT down.

HEA-NOC/20080116-2 [CLOSE] Link to GEANT down.

Gareth Eason <[log in to unmask]>

Thu, 24 Jan 2008 12:44:36 +0000

49 lines

New Thread

HEA-NOC/20080116-2 [OPEN] Link to GEANT down.

HEA-NOC/20080116-2 [OPEN] Link to GEANT down.

Gareth Eason <[log in to unmask]>

Wed, 16 Jan 2008 07:31:15 +0000

28 lines

New Thread

HEA-NOC/20080116-2 [UPDATE] Link to GEANT down.

HEA-NOC/20080116-2 [UPDATE] Link to GEANT down.

Gareth Eason <[log in to unmask]>

Wed, 23 Jan 2008 10:22:23 +0000

44 lines

HEA-NOC/20080116-2 [UPDATE] Link to GEANT down.

Justin Hourigan <[log in to unmask]>

Thu, 17 Jan 2008 09:07:06 +0000

41 lines

HEA-NOC/20080116-2 [UPDATE] Link to GEANT down.

Gareth Eason <[log in to unmask]>

Wed, 16 Jan 2008 08:14:22 +0000

35 lines

New Thread

HEA-NOC/20080116-7 [CLOSE] BT fibre maintanence on Cork-Limerick GigE link

HEA-NOC/20080116-7 [CLOSE] BT fibre maintanence on Cork-Limerick GigE link

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 16:22:40 +0000

43 lines

New Thread

HEA-NOC/20080116-7 [OPEN] BT fibre maintanence on Cork-Limerick GigE link

HEA-NOC/20080116-7 [OPEN] BT fibre maintanence on Cork-Limerick GigE link

Justin Hourigan <[log in to unmask]>

Wed, 16 Jan 2008 11:12:44 +0000

39 lines

New Thread

HEA-NOC/20080118-3 [OPEN] Both circuits to the IEDR are reporting down

HEA-NOC/20080118-3 [OPEN] Both circuits to the IEDR are reporting down

Brian Boyle <[log in to unmask]>

Fri, 18 Jan 2008 14:36:06 +0000

25 lines

New Thread

HEA-NOC/20080118-3 [UPDATE] Both circuits to the IEDR are reporting down

HEA-NOC/20080118-3 [UPDATE] Both circuits to the IEDR are reporting down

Justin Hourigan <[log in to unmask]>

Fri, 18 Jan 2008 14:39:37 +0000

33 lines

New Thread

HEA-NOC/20080118-4 [CLOSE] UL website not responding

HEA-NOC/20080118-4 [CLOSE] UL website not responding

Justin Hourigan <[log in to unmask]>

Fri, 18 Jan 2008 15:57:43 +0000

37 lines

New Thread

HEA-NOC/20080118-4 [OPEN] UL website not responding

HEA-NOC/20080118-4 [OPEN] UL website not responding

Justin Hourigan <[log in to unmask]>

Fri, 18 Jan 2008 15:08:57 +0000

27 lines

New Thread

HEA-NOC/20080118-4 [UPDATE] UL website not responding

HEA-NOC/20080118-4 [UPDATE] UL website not responding

Justin Hourigan <[log in to unmask]>

Fri, 18 Jan 2008 15:26:31 +0000

34 lines

New Thread

HEA-NOC/20080119-2 [CLOSE] GMIT inaccessible

HEA-NOC/20080119-2 [CLOSE] GMIT inaccessible

Dave Wilson <[log in to unmask]>

Mon, 21 Jan 2008 10:23:22 +0000

41 lines

New Thread

HEA-NOC/20080119-2 [OPEN] GMIT inaccessible

HEA-NOC/20080119-2 [OPEN] GMIT inaccessible

Dave Wilson <[log in to unmask]>

Sat, 19 Jan 2008 19:25:24 +0000

32 lines

New Thread

HEA-NOC/20080119-2 [UPDATE] GMIT inaccessible

HEA-NOC/20080119-2 [UPDATE] GMIT inaccessible

Dave Wilson <[log in to unmask]>

Sun, 20 Jan 2008 09:41:22 +0000

37 lines

New Thread

HEA-NOC/20080121-12 [OPEN] ESBT Planned Outage Notice

HEA-NOC/20080121-12 [OPEN] ESBT Planned Outage Notice

Cormac O'Cianain <[log in to unmask]>

Mon, 21 Jan 2008 16:56:41 +0000

29 lines

New Thread

HEA-NOC/20080121-2 [CLOSE] ESB Outage in Dundalk

HEA-NOC/20080121-2 [CLOSE] ESB Outage in Dundalk

Cormac O'Cianain <[log in to unmask]>

Mon, 21 Jan 2008 15:01:30 +0000

38 lines

New Thread

HEA-NOC/20080121-2 [OPEN] ESB Outage in Dundalk

HEA-NOC/20080121-2 [OPEN] ESB Outage in Dundalk

Cormac O'Cianain <[log in to unmask]>

Mon, 21 Jan 2008 11:22:52 +0000

29 lines

New Thread

HEA-NOC/20080121-2 [UPDATE] ESB Outage in Dundalk

HEA-NOC/20080121-2 [UPDATE] ESB Outage in Dundalk

Cormac O'Cianain <[log in to unmask]>

Mon, 21 Jan 2008 12:04:21 +0000

33 lines

New Thread

HEA-NOC/20080121-9 [OPEN] IOS Upgrades

HEA-NOC/20080121-9 [OPEN] IOS Upgrades

Garwin Liu <[log in to unmask]>

Mon, 21 Jan 2008 15:45:53 +0000

31 lines

New Thread

HEA-NOC/20080122-5 [CLOSE] Maintenance Work CDVEC

HEA-NOC/20080122-5 [CLOSE] Maintenance Work CDVEC

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 17:23:35 +0000

35 lines

New Thread

HEA-NOC/20080122-5 [OPEN] Maintenance Work CDVEC

HEA-NOC/20080122-5 [OPEN] Maintenance Work CDVEC

Cormac O'Cianain <[log in to unmask]>

Tue, 22 Jan 2008 12:18:26 +0000

29 lines

New Thread

HEA-NOC/20080122-6 [CLOSE] DIT migration to new IP equipment

HEA-NOC/20080122-6 [CLOSE] DIT migration to new IP equipment

Daniel Lete <[log in to unmask]>

Wed, 23 Jan 2008 08:19:30 +0000

48 lines

New Thread

HEA-NOC/20080122-6 [OPEN] DIT migration to new IP equipment

HEA-NOC/20080122-6 [OPEN] DIT migration to new IP equipment

Daniel Lete <[log in to unmask]>

Tue, 22 Jan 2008 13:03:32 +0000

44 lines

New Thread

HEA-NOC/20080122-9 [CLOSE] EPA Firewall Testing

HEA-NOC/20080122-9 [CLOSE] EPA Firewall Testing

Cormac O'Cianain <[log in to unmask]>

Mon, 28 Jan 2008 11:39:17 +0000

44 lines

New Thread

HEA-NOC/20080122-9 [OPEN] EPA Firewall TestingThe

HEA-NOC/20080122-9 [OPEN] EPA Firewall TestingThe

Cormac O'Cianain <[log in to unmask]>

Tue, 22 Jan 2008 15:06:50 +0000

28 lines

New Thread

HEA-NOC/20080122-9 [UPDATE] EPA Firewall Testing

HEA-NOC/20080122-9 [UPDATE] EPA Firewall Testing

Cormac O'Cianain <[log in to unmask]>

Mon, 28 Jan 2008 10:53:25 +0000

38 lines

HEA-NOC/20080122-9 [UPDATE] EPA Firewall Testing

Cormac O'Cianain <[log in to unmask]>

Thu, 24 Jan 2008 15:32:32 +0000

36 lines

New Thread

HEA-NOC/20080122-9 [UPDATE] EPA Firewall TestingThe

HEA-NOC/20080122-9 [UPDATE] EPA Firewall TestingThe

Cormac O'Cianain <[log in to unmask]>

Tue, 22 Jan 2008 16:15:06 +0000

33 lines

New Thread

HEA-NOC/20080123-10 [CLOSE] mirror services degraded.

HEA-NOC/20080123-10 [CLOSE] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Wed, 30 Jan 2008 14:56:19 +0000

47 lines

New Thread

HEA-NOC/20080123-10 [OPEN] mirror services degraded.

HEA-NOC/20080123-10 [OPEN] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Wed, 23 Jan 2008 13:30:28 +0000

29 lines

New Thread

HEA-NOC/20080123-10 [UPDATE] mirror services degraded.

HEA-NOC/20080123-10 [UPDATE] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Mon, 28 Jan 2008 16:07:13 +0000

42 lines

HEA-NOC/20080123-10 [UPDATE] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Mon, 28 Jan 2008 12:20:35 +0000

40 lines

HEA-NOC/20080123-10 [UPDATE] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Mon, 28 Jan 2008 09:53:33 +0000

36 lines

HEA-NOC/20080123-10 [UPDATE] mirror services degraded.

Brian Scanlan <[log in to unmask]>

Mon, 28 Jan 2008 09:10:59 +0000

33 lines

New Thread

HEA-NOC/20080123-11 [CLOSE] Migration of LISTSERV to new server.

HEA-NOC/20080123-11 [CLOSE] Migration of LISTSERV to new server.

Brian Scanlan <[log in to unmask]>

Wed, 30 Jan 2008 14:55:38 +0000

41 lines

New Thread

HEA-NOC/20080123-11 [OPEN] Migration of LISTSERV to new server.

HEA-NOC/20080123-11 [OPEN] Migration of LISTSERV to new server.

Brian Scanlan <[log in to unmask]>

Wed, 23 Jan 2008 14:14:44 +0000

28 lines

New Thread

HEA-NOC/20080123-11 [UPDATE] Migration of LISTSERV to new server.

HEA-NOC/20080123-11 [UPDATE] Migration of LISTSERV to new server.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 16:51:43 +0000

36 lines

HEA-NOC/20080123-11 [UPDATE] Migration of LISTSERV to new server.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 08:59:56 +0000

31 lines

HEA-NOC/20080123-11 [UPDATE] Migration of LISTSERV to new server.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 09:16:23 +0000

33 lines

New Thread

HEA-NOC/20080123-12 [CLOSE] Change of Bluenet PE node interfaces from routing to switching

HEA-NOC/20080123-12 [CLOSE] Change of Bluenet PE node interfaces from routing to switching

Gareth Eason <[log in to unmask]>

Thu, 31 Jan 2008 10:04:10 +0000

91 lines

New Thread

HEA-NOC/20080123-12 [OPEN] Change of Bluenet PE node interfaces from routing to switching

HEA-NOC/20080123-12 [OPEN] Change of Bluenet PE node interfaces from routing to switching

Gareth Eason <[log in to unmask]>

Wed, 23 Jan 2008 18:01:55 +0000

76 lines

New Thread

HEA-NOC/20080123-12 [UPDATE] Change of Bluenet PE node interfaces from routing to switching

HEA-NOC/20080123-12 [UPDATE] Change of Bluenet PE node interfaces from routing to switching

Gareth Eason <[log in to unmask]>

Wed, 30 Jan 2008 08:37:00 +0000

88 lines

HEA-NOC/20080123-12 [UPDATE] Change of Bluenet PE node interfaces from routing to switching

Gareth Eason <[log in to unmask]>

Wed, 30 Jan 2008 07:06:40 +0000

82 lines

HEA-NOC/20080123-12 [UPDATE] Change of Bluenet PE node interfaces from routing to switching

Louis Twomey <[log in to unmask]>

Tue, 29 Jan 2008 10:12:55 +0000

79 lines

New Thread

HEA-NOC/20080123-8 [CLOSE] Loss of connectivity

HEA-NOC/20080123-8 [CLOSE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Thu, 24 Jan 2008 15:48:41 +0000

35 lines

New Thread

HEA-NOC/20080123-8 [OPEN] Loss of connectivity

HEA-NOC/20080123-8 [OPEN] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Wed, 23 Jan 2008 12:11:36 +0000

24 lines

New Thread

HEA-NOC/20080123-8 [UPDATE] Loss of connectivity

HEA-NOC/20080123-8 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Wed, 23 Jan 2008 13:36:35 +0000

30 lines

HEA-NOC/20080123-8 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Wed, 23 Jan 2008 12:27:02 +0000

27 lines

New Thread

HEA-NOC/20080124-1 [OPEN] DIAS IOS upgrade.

HEA-NOC/20080124-1 [OPEN] DIAS IOS upgrade.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 08:50:51 +0000

28 lines

New Thread

HEA-NOC/20080124-1 [UPDATE] DIAS IOS upgrade.

HEA-NOC/20080124-1 [UPDATE] DIAS IOS upgrade.

Brian Scanlan <[log in to unmask]>

Thu, 24 Jan 2008 08:55:41 +0000

32 lines

New Thread

HEA-NOC/20080124-4 [OPEN] Janet connection down

HEA-NOC/20080124-4 [OPEN] Janet connection down

Cormac O'Cianain <[log in to unmask]>

Thu, 24 Jan 2008 11:22:51 +0000

28 lines

New Thread

HEA-NOC/20080124-4 [UPDATE] Janet connection down

HEA-NOC/20080124-4 [UPDATE] Janet connection down

Gareth Eason <[log in to unmask]>

Tue, 29 Jan 2008 10:09:07 +0000

37 lines

New Thread

HEA-NOC/20080124-5 [CLOSE] Loss of connectivity Outreach Centre UL

HEA-NOC/20080124-5 [CLOSE] Loss of connectivity Outreach Centre UL

Cormac O'Cianain <[log in to unmask]>

Thu, 24 Jan 2008 12:50:07 +0000

35 lines

New Thread

HEA-NOC/20080124-5 [OPEN] Loss of connectivity Outreach Centre UL

HEA-NOC/20080124-5 [OPEN] Loss of connectivity Outreach Centre UL

Cormac O'Cianain <[log in to unmask]>

Thu, 24 Jan 2008 12:02:08 +0000

28 lines

New Thread

HEA-NOC/20080125-1 [CLOSE] Loss of connectivity

HEA-NOC/20080125-1 [CLOSE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 12:39:22 +0000

43 lines

New Thread

HEA-NOC/20080125-1 [OPEN] Loss of connectivity

HEA-NOC/20080125-1 [OPEN] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 09:31:01 +0000

26 lines

New Thread

HEA-NOC/20080125-1 [UPDATE] Loss of connectivity

HEA-NOC/20080125-1 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 11:21:09 +0000

38 lines

HEA-NOC/20080125-1 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 11:09:36 +0000

35 lines

HEA-NOC/20080125-1 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 10:22:55 +0000

32 lines

HEA-NOC/20080125-1 [UPDATE] Loss of connectivity

Cormac O'Cianain <[log in to unmask]>

Fri, 25 Jan 2008 09:43:24 +0000

29 lines

New Thread

HEA-NOC/20080125-3 [CLOSE] Upgrade coupland to FreeBSD 6.3

HEA-NOC/20080125-3 [CLOSE] Upgrade coupland to FreeBSD 6.3

Rob Gallagher <[log in to unmask]>

Fri, 25 Jan 2008 11:58:21 +0000

34 lines

New Thread

HEA-NOC/20080125-3 [OPEN] Upgrade coupland to FreeBSD 6.3

HEA-NOC/20080125-3 [OPEN] Upgrade coupland to FreeBSD 6.3

Rob Gallagher <[log in to unmask]>

Fri, 25 Jan 2008 10:33:09 +0000

28 lines

New Thread

HEA-NOC/20080126-1 [CLOSE] Hermes intermittent outages

HEA-NOC/20080126-1 [CLOSE] Hermes intermittent outages

Louis Twomey <[log in to unmask]>

Thu, 31 Jan 2008 17:47:53 +0000

41 lines

New Thread

HEA-NOC/20080126-1 [OPEN] Hermes intermittent outages

HEA-NOC/20080126-1 [OPEN] Hermes intermittent outages

Cillian Sharkey <[log in to unmask]>

Sat, 26 Jan 2008 11:45:22 +0000

28 lines

New Thread

HEA-NOC/20080126-1 [UPDATE] Hermes intermittent outages

HEA-NOC/20080126-1 [UPDATE] Hermes intermittent outages

Louis Twomey <[log in to unmask]>

Mon, 28 Jan 2008 13:00:20 +0000

35 lines

HEA-NOC/20080126-1 [UPDATE] Hermes intermittent outages

Louis Twomey <[log in to unmask]>

Mon, 28 Jan 2008 10:39:29 +0000

32 lines

New Thread

HEA-NOC/20080126-5 [OPEN] Bluenet Dublin ring brief loss of resilience

HEA-NOC/20080126-5 [OPEN] Bluenet Dublin ring brief loss of resilience

Cillian Sharkey <[log in to unmask]>

Sat, 26 Jan 2008 12:56:05 +0000

31 lines

New Thread

HEA-NOC/20080127-1 [CLOSE] GEANT primary circuit outage

HEA-NOC/20080127-1 [CLOSE] GEANT primary circuit outage

Louis Twomey <[log in to unmask]>

Mon, 28 Jan 2008 10:36:08 +0000

38 lines

New Thread

HEA-NOC/20080127-1 [OPEN] GEANT primary circuit outage

HEA-NOC/20080127-1 [OPEN] GEANT primary circuit outage

Cillian Sharkey <[log in to unmask]>

Sun, 27 Jan 2008 11:05:35 +0000

27 lines

New Thread

HEA-NOC/20080127-1 [UPDATE] GEANT primary circuit outage

HEA-NOC/20080127-1 [UPDATE] GEANT primary circuit outage

Cillian Sharkey <[log in to unmask]>

Sun, 27 Jan 2008 15:21:53 +0000

31 lines

New Thread

HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap

HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap

Louis Twomey <[log in to unmask]>

Mon, 28 Jan 2008 15:39:24 +0000

37 lines

New Thread

HEA-NOC/20080127-5 [OPEN] St. Angela's brief flap

HEA-NOC/20080127-5 [OPEN] St. Angela's brief flap

Cillian Sharkey <[log in to unmask]>

Sun, 27 Jan 2008 16:02:46 +0000

28 lines

New Thread

HEA-NOC/20080127-5 [REOPEN] St. Angela's brief flap

HEA-NOC/20080127-5 [REOPEN] St. Angela's brief flap

Louis Twomey <[log in to unmask]>

Thu, 31 Jan 2008 09:59:28 +0000

42 lines

New Thread

HEA-NOC/20080127-5 [UPDATE] St. Angela's brief flap

HEA-NOC/20080127-5 [UPDATE] St. Angela's brief flap

Louis Twomey <[log in to unmask]>

Thu, 31 Jan 2008 16:49:55 +0000

49 lines

HEA-NOC/20080127-5 [UPDATE] St. Angela's brief flap

Louis Twomey <[log in to unmask]>

Thu, 31 Jan 2008 10:17:07 +0000

44 lines

HEA-NOC/20080127-5 [UPDATE] St. Angela's brief flap

Cillian Sharkey <[log in to unmask]>

Sun, 27 Jan 2008 17:54:21 +0000

32 lines

New Thread

HEA-NOC/20080128-2 [CLOSE] IT Tallaght migration to new IP equipment

HEA-NOC/20080128-2 [CLOSE] IT Tallaght migration to new IP equipment

Daniel Lete <[log in to unmask]>

Mon, 28 Jan 2008 15:02:31 +0000

50 lines

New Thread

HEA-NOC/20080128-2 [UPDATE] IT Tallaght migration to new IP equipment

HEA-NOC/20080128-2 [UPDATE] IT Tallaght migration to new IP equipment

Daniel Lete <[log in to unmask]>

Mon, 28 Jan 2008 14:27:44 +0000

43 lines

New Thread

HEA-NOC/20080129-10 [OPEN] Upgrade to SNMP module on centralised DC Rectifier

HEA-NOC/20080129-10 [OPEN] Upgrade to SNMP module on centralised DC Rectifier

Gareth Eason <[log in to unmask]>

Tue, 29 Jan 2008 17:55:02 +0000

38 lines

New Thread

HEA-NOC/20080129-2 [CLOSE] Outage on primary GEANT link

HEA-NOC/20080129-2 [CLOSE] Outage on primary GEANT link

Louis Twomey <[log in to unmask]>

Wed, 30 Jan 2008 11:56:53 +0000

35 lines

New Thread

HEA-NOC/20080129-2 [OPEN] Outage on primary GEANT link

HEA-NOC/20080129-2 [OPEN] Outage on primary GEANT link

Louis Twomey <[log in to unmask]>

Tue, 29 Jan 2008 10:07:25 +0000

30 lines

New Thread

HEA-NOC/20080129-9 [OPEN] WIT migration to new IP equipment

HEA-NOC/20080129-9 [OPEN] WIT migration to new IP equipment

Cormac O'Cianain <[log in to unmask]>

Tue, 29 Jan 2008 17:16:56 +0000

40 lines

New Thread

HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment

HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment

Cormac O'Cianain <[log in to unmask]>

Thu, 31 Jan 2008 09:16:10 +0000

44 lines

New Thread

HEA-NOC/20080130-2 [CLOSE] Loss of connectivity to parts of Services network in Blanch

HEA-NOC/20080130-2 [CLOSE] Loss of connectivity to parts of Services network in Blanch

Louis Twomey <[log in to unmask]>

Thu, 31 Jan 2008 17:46:54 +0000

51 lines

New Thread

HEA-NOC/20080130-2 [OPEN] Loss of connectivity to parts of Services network in Blanch

HEA-NOC/20080130-2 [OPEN] Loss of connectivity to parts of Services network in Blanch

Louis Twomey <[log in to unmask]>

Wed, 30 Jan 2008 10:39:15 +0000

39 lines

New Thread

HEA-NOC/20080130-2 [UPDATE] Loss of connectivity to parts of Services network in Blanch

HEA-NOC/20080130-2 [UPDATE] Loss of connectivity to parts of Services network in Blanch

Louis Twomey <[log in to unmask]>

Wed, 30 Jan 2008 13:26:03 +0000

45 lines

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

July 2017
June 2017
May 2017
April 2017
March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager